On a 40-km bike ride, a cyclist rides the first 20 km at 20 km/h.If the cyclist rides very fast for the final 20 km, what is the maximum value his average speed could approach?


1
Expert's answer
2021-12-02T13:58:40-0500

Explanations & Calculations


  • Write a relationship for the average speed and integrate it to see the maximum.

"\qquad\qquad
\begin{aligned}
\small V_{avg}&=\small \frac{40\,km}{\frac{20\,km}{20\,kmh^{-1}}+\frac{20\,km}{v}}\\
&=\small \frac{40}{\Big(1+\frac{20}{v}\Big)}
\end{aligned}"

  • For the average speed to be maximum, the denominator needs to be minimum which is "\small 1" .
  • For any speed during the final 20 kilometres, the denominator value is beyond 1.
  • Therefore, the maximum average speed he could attain is "\small 40\,kmh^{-1}" .